OpenStudy (mathmath333):

the sum of the cubes of the digits constituting a two digit number is 243 and the product of the sum of its digit by the product of its digit is 162.find the two digit number.

OpenStudy (anonymous):

half of this is somewhere on yahoo im telling ya

OpenStudy (mathmath333):

i got it

OpenStudy (mathmath333):

a+c=9

OpenStudy (mathmath333):

not from yahoo though

OpenStudy (anonymous):

product of the sum of its digit by the product of its digit is this is confusing

ganeshie8 (ganeshie8):

whats ur solution ?

OpenStudy (mathmath333):

wait 1 min

OpenStudy (anonymous):

So does that mean (x+y)/xy = 162

OpenStudy (mathmath333):

63,36

OpenStudy (mathmath333):

no 

OpenStudy (mathmath333):

(x+y)/xy = 162 this will spin ur head

OpenStudy (anonymous):

and first one is (10x)^3 +(y)^3 right

OpenStudy (anonymous):

just asking , because i am trying to solve the problem too

OpenStudy (mathmath333):